A marble is thrown horizontally from a tabletop with a velocity of 1.50 m/s. The marble falls 0.70 m away from the table’s edge. How high is the lab table? What is the marble’s velocity just before it hits the floor? 


1
Expert's answer
2021-06-02T13:12:09-0400

Gives

Velocity (v0x)=1.50m/sec

d=0.70m

We know that

"d=V_{0x}t"

"0.70=1.50\times t"

t=0.47 sec

"V_{fy}=V_{0y}+gt"

"V_{fy}=0+(-9.8)\times0.47"

"V_{fy}=-4.5783m\/sec"

"V=\sqrt{V_{ox}^2+V_{fy}^2}"

Put value

"V=\sqrt{{1.50}^2+{(-0.47)}^2}"

V=4.82m/sec